A saline solution is prepared by dissolving 5.04 g of NaCl in 95.0 g of water.



Calculate the following:



a. Mole fraction of the solute and the solvent



b. % mass of the solute



c. Concentration of solution in ppm




Please put what letter did you answer in every solution

1
Expert's answer
2022-03-10T07:30:05-0500

Solution:

solute = NaCl

solvent = H2O


The molar mass of NaCl is 58.44 g/mol

The molar mass of water (H2O) is 18.015 g/mol


Calculate the moles of each component:

Moles of NaCl = (5.04 g NaCl) × (1 mol NaCl / 58.44 g NaCl) = 0.0862 mol NaCl

Moles of H2O = (95.0 g H2O) × (1 mol H2O / 18.015 g H2O) = 5.2734 mol H2O

Therefore,

Total moles of solution = Moles of NaCl + Moles of H2O = 0.0862 mol + 5.2734 mol = 5.3596 mol


Calculate the mole fraction (χ) of each component:

χ = Moles of component / Total moles of solution

Therefore,

χsolute = χNaCl = Moles of NaCl / Total moles of solution = (0.0862 mol) / (5.3596 mol) = 0.016

χsolute = 0.016

χsolvent = χH2O = Moles of H2O / Total moles of solution = (5.2734 mol) / (5.3596 mol) = 0.984

χsolvent = 0.984


Calculate the %mass of each component:

Percent by mass = (Mass of component / Mass of solution) × 100%

Mass of solution = Mass of NaCl + Mass of H2O = 5.04 g  + 95.0 g = 100.04 g

Therefore,

%mass of solute = %mass of NaCl = (5.04 g / 100.04g) × 100% = 5.04%

%mass of solute = 5.04%

%mass of solvent = %mass of H2O = (95.0 g / 100.04g) × 100% = 94.96%

%mass of solvent = 94.96%


Calculate the concentration of solution in ppm:

ppm = parts per million = (grams of solute / grams of solution) × 1000000

Therefore,

Concentration of solution = (Mass of NaCl / Mass of solution) × 1000000

Concentration of solution = (5.04 g / 100.04g) × 1000000 = 50379.85 ppm = 50380 ppm

Concentration of solution = 50380 ppm


Answers:

a)

The mole fraction of the solute (χsolute) is 0.016

The mole fraction of the solvent (χsolvent) is 0.984

b)

The percent by mass of the solute (%mass of solute) is 5.04%

c)

The concentration of solution is 50380 ppm